The median existing home price of a home is 1988 was $90,600.00, and in 1994 it was $107,600.00. You wish to fit a linear equation to the points representing this data using 1980 as the base year.
a.​Write and solve a system of equations that fit this model.
b.​Interpret the slope in the context of the data.​
c.​By your model, what was the median price in 1991?

Answer by Boreal(15235)

price P=(year-1980)x+ some constant
ordered pair (8, 90600) and (14, 107,600)
slope is difference in y 17000/ difference in x 6,
that is $2833.3
point slope formula is y-y1=m(x-x1), m slope and (x1, y1) a point
y-90600=2833.3(x-8)
y=2833.3x+67934
slope is increase in the price per year.
1991 is x=11
so price was in 1991: 2833.33*11+67934=$99,100.67